OpenStudy (arihangdu):

Please! Please can anybody help me this problem? A sound wave is modeled with the equation y = 1/4 cos 2pi/3 θ. a. Find the period. Explain your method. b. Find the amplitude. Explain your method. C. What is the equation of the midline? What does it represent? Please solve all will get madel and fan.

OpenStudy (md152727):

note that waves can be of general eq, y=r cos wt, where r is the amplitude, w the angular velocity and t is the time parameter.

OpenStudy (md152727):

cross reference your eq with the general eq above.

OpenStudy (arihangdu):

Ok

OpenStudy (md152727):

ω=2πT

OpenStudy (arihangdu):

Actually this is a algebra problem

OpenStudy (md152727):

i know

OpenStudy (md152727):

amplitude, A through a mathematical analysis, the amplitude is the highest/lowest point of the curve. y=14cos2π3θ y′=−14(2π3)sin2π3θ let y'=0, sin2π3θ=0 θ=0 y=1/4=A you follow?

OpenStudy (md152727):

now, y′=ω the period is the time for two wavecrest, i.e. two amplitude, also aka two y'=0 so you find the w for two successive max points and you'll get it.

OpenStudy (arihangdu):

OK I'm following you

OpenStudy (arihangdu):

I didn't understand please help me to solve all.

OpenStudy (arihangdu):

Can anybody do this please please please help the friends I need.

OpenStudy (arihangdu):

@MD152727 please please help me to solve this problem with explain

OpenStudy (anonymous):

A

Directrix (directrix):

Look at this thread for help. You'll need to read the whole thing. http://openstudy.com/study#/updates/519e5489e4b04449b22151a5 @Arihangdu

OpenStudy (arihangdu):

Yes I know to find a and b now but what is for C?

OpenStudy (arihangdu):

@Directrix what is the equation of the midline? What does it represent? Please

Directrix (directrix):

The midline is a horizontal axis that is used as the reference line about which the graph of a trigonometric function oscillates. The equation of the midline of trignometric function is the average of the maximum and minimum values of the function i.e. y=(min. + max.)/2.

Directrix (directrix):

You might find this information helpful: http://www.enotes.com/homework-help/how-midline-used-help-graph-trigonometric-465053

OpenStudy (arihangdu):

Ufff I never get the midline hat maths

OpenStudy (arihangdu):

Please how to find min+max?..

OpenStudy (arihangdu):

You said (min+ max)/2 But you did (min- max)/2. ?

OpenStudy (arihangdu):

I found amplitude =. -1/4 Period = 3 Is that right?

Directrix (directrix):

You said (min+ max)/2 But you did (min- max)/2. 1/4 + (-1/4) = 1/4 - 1/4 = 0 y = 0.

Directrix (directrix):

>>>I found amplitude =. -1/4 No. I got +1/4 for the amp and period 3.

OpenStudy (arihangdu):

Thank you for helping dear. Appreciate your help thanks.